# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 7.2. Generic KSBA stability of the mirror family
To prove ii), it is enough to prove that $Y_{t}$ is normal, since then the other conditions follow by i). By Theorem 7.5, the pair $(Y_{t},D_{t})$ is semi-log-canonical. In particular, the $Y_{t}$ is $S_{2}$, and so, by Serre’s criterion for normality [51, Thm 8.22A], it suffices to show that $Y_{t}$ is regular in codimension one. As in the proofs of Lemma 7.3 and Theorem 7.5, we can assume that there exists a good divisor $F$ on $\mathcal{X}$, inducing a $\mathbb{C}^{\star}$-action on $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$. The closure of the $\mathbb{C}^{\star}$-orbit passing through $t$ is an affine line $\mathbb{A}^{1}$, with $(\mathbb{A}^{1}\setminus\{0\})\subset T_{\mathcal{X}}$, and $0\in\mathrm{Spec}\,\mathbb{C}[G]\simeq\mathrm{Spec}\,\mathbb{C}[H]/J$, where $G=F^{\perp}\cap NE(\mathcal{X}/\overline{\mathcal{X}})$ and $J=\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]\setminus J$. The restriction $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$ of $\mathcal{Y}_{\mathcal{X}}$ to this affine line is constant with fibers $Y_{t}$ on $\mathbb{A}^{1}\setminus\{0\}$. Moreover, by the proof of Lemma 7.2, the central fiber $\mathcal{Y}_{\mathcal{X},0}$ of $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$ is a union of toric varieties with intersection complex $(P,\mathscr{P}_{J})$, where $\mathscr{P}_{J}$ is the subdivision of $P$ obtained from $\mathscr{P}_{J}$ by omitting the edges $e$ such that the corresponding curves $C_{e}$ in the double locus of $\mathcal{X}_{0}$ satisfy $F\cdot C_{e}=0$. In particular, the singular locus of $(P,\mathscr{P}_{J})$ is the union of its double curves $D_{e}$, which are in one-to-one correspondence with the edges $e$ of $\mathscr{P}_{J}$, that is, with the double curves $C_{e}$ of $\mathcal{X}_{0}$ such that $F\cdot C_{e}\neq 0$. Hence to prove that $Y_{t}$ is smooth in codimension one, it is enough to show that each of the double curves $D_{e}$ are generically smoothed out in $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$.
@@
It follows from the construction of $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ from the canonical scattering diagram, together with [45, Eq.(2.17)], that a neighborhood of $D_{e}$ in $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$ is isomorphic to the affine variety of equation $xy=t^{[C_{e}]}f_{W,e}$, where $f_{W,e}$ is the function of the canonical scattering diagram attached to the wall supported on the edge $e$ of $P$. As $(\mathbb{A}^{1}\setminus\{0\})$ is contained in $T_{\mathcal{X}}$, we have $t^{[C_{e}]}\neq 0$ on the non-central fibers of $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$. Moreover, by Lemma 6.9, the function $f_{W,e}$ is generically non-zero on $D_{e}$. This shows that all the double curves $D_{e}$ of $\mathcal{Y}_{\mathcal{X},0}$ are generically smoothed out in the one-parameter family $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$, and this concludes the proof that $Y_{t}$ is normal. ∎
@@ 7.3. KSBA stability of the mirror family
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, and let $H$ be a face of $NE(\mathcal{X}/\overline{\mathcal{X}})$. As $f:\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is a relative Mori dream space by Proposition 3.36, there is a corresponding birational contraction $f_{H}:\mathcal{X}\rightarrow\mathcal{X}_{H}$ such that $H=NE(\mathcal{X}/\mathcal{X}_{H})$. Denote by $S_{H}:=\mathrm{Spec}\,\mathbb{C}[H]=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\mathcal{X}_{H})]$ the corresponding stratum of $S_{\mathcal{X}}$, with dense torus orbit $T_{H}:=\mathrm{Spec}\,\mathbb{C}[H^{\mathrm{gp}}]$. Let $\mathscr{P}_{H}$ be the polyhedral decomposition of $P$ obtained as a coarsening of $\mathscr{P}$ by removing all edges $e$ of $\mathscr{P}$ such that the corresponding double curves $C_{e}$ of $\mathcal{X}_{0}$ satisfy $[C_{e}]\in H$, that is, are contracted by $f_{H}$. For every 2-dimensional cell $\sigma$ of $\mathscr{P}_{H}$, we denote by $\mathscr{P}_{\sigma}$ the polyhedral decomposition of $\sigma$ obtained by restricting $\mathscr{P}$ to $\sigma$.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration and let $H$ be a face of $NE(\mathcal{X}/\overline{\mathcal{X}})$. Then, for every 2-dimensional cell $\sigma$ of $\mathscr{P}_{H}$, there exists a quasi-projective open Kulikov surface $\mathcal{X}^{\prime}_{\sigma,0}$, with dual intersection complex $(\sigma,\mathscr{P}_{\sigma})$, such that $\mathcal{X}_{\sigma,0}$ is a Zariski open dense subset of $\mathcal{X}^{\prime}_{\sigma,0}$, containing all compact curves of $\mathcal{X}^{\prime}_{\sigma,0}$, and $U_{\sigma}$ is a Zariski open neighborhood of the central point in the affinization $\overline{\mathcal{X}}^{\prime}_{\sigma,0}$.
@@
First note that $\mathcal{X}_{\sigma,0}$ is quasi-projective, reduced and normal crossings, as an open subset of the quasi-projective, reduced and normal crossings surface $\mathcal{X}_{0}$. If $v\in\sigma_{\mathbb{Z}}$ and $v\notin\partial\sigma$, then it follows from Lemma 7.8 ii) that $X^{v}_{\sigma}=X^{v}$, and so $X^{v}_{\sigma}$ is a 2-surface. If $v\in\partial\sigma$, then it follows from Definitions 3.2-3.4 that $X^{v}_{\sigma}$ is naturally a Zariski open subset of a 0-surface or a 1-surface if $\dim f_{H}(X^{v}_{\sigma})=2$ or $\dim f_{H}(X^{v}_{\sigma})=1$ respectively. Finally, it follows from [86, Proposition 3.3] that $\sigma$ is homeomorphic to a disk. ∎
Let $\mathcal{X}^{\prime}_{\sigma,0}$ be the quasi-projective open Kulikov surface given by Lemma 7.9. By Theorem 3.21, we can assume up to locally trivial deformations that $\mathcal{X}_{\sigma,0}^{\prime}$ is generic $d$-semistable, and so that $\mathcal{X}_{\sigma,0}^{\prime}$ is the central fiber of a quasi-projective open Kulikov degeneration $\mathcal{X}_{\sigma}\rightarrow\Delta$ by Theorem 3.26 and Lemma 3.31. The inclusions $\mathcal{X}_{\sigma,0}\subset\mathcal{X}_{0}$ and $\overline{\mathcal{X}}_{\sigma,0}\subset\mathcal{X}_{H,0}$ induce an injective map $NE(\mathcal{X}_{\sigma}/\overline{\mathcal{X}}_{\sigma})\rightarrow H=NE(\mathcal{X}/\mathcal{X}_{H})$, and so a dominant toric morphism
$S_{H}=\mathrm{Spec}\,\mathbb{C}[H]\longrightarrow S_{\mathcal{X}_{\sigma}}=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\sigma}/\overline{\mathcal{X}}_{\sigma})]\,.$ (7.1)
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, and let $H$ be a face of $NE(\mathcal{X}/\overline{\mathcal{X}})$. Then, for every $\beta\in H$ and $p,q,r\in C(P)_{\mathbb{Z}}$, we have $N_{pqr}^{\beta}(\mathcal{X})=0$ unless $p,q,r$ are all contained in a common cone $C(\sigma)$ over a 2-dimensional cell $\sigma$ of $\mathscr{P}_{H}$.
@@
The tropicalization of a punctured log curve $C$ contributing to $N_{pqr}^{\beta}(\mathcal{X})$ as in Definition 5.3 is a connected tropical curve in $C(P)$, with two unbounded edges with asymptotic directions $\mathbb{R}_{\geq 0}p$ and $\mathbb{R}_{\geq 0}q$, and with one leg contained in $\mathbb{R}_{\geq 0}r$. In particular, if $p,q,r$ were not all contained in a common cone $C(\sigma)$, this tropical curve would have to intersect a cone $C(\sigma\cap\sigma^{\prime})$ over the intersection of two distinct 2-dimensional faces $\sigma$, $\sigma^{\prime}$ of $\mathscr{P}_{\sigma}$, and so a cone $C(e)$ over an edge $e$ of $\mathscr{P}_{\sigma}$. In particular, the curve $C$ would contain an irreducible component with class a multiple of $[C_{e}]$, where $C_{e}$ is the double curve of $\mathcal{X}_{0}$ corresponding to $e$. Since $e$ is an edge of $\mathscr{P}_{\sigma}$, we have $[C_{e}]\notin H$, and this would imply $\beta\notin H$. Hence, the results follows since we assume that $\beta\in H$. ∎
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration and let $H$ be a face of $NE(\mathcal{X}/\overline{\mathcal{X}})$. Then, for every 2-dimensional cell $\sigma$ of $\mathscr{P}_{H}$, for every $p,q,r\in C(\sigma)_{\mathbb{Z}}$ and $\beta\in H$, we have $N_{pqr}^{\beta}(\mathcal{X})=0$ if $\beta\notin NE(\mathcal{X}_{\sigma}/\overline{\mathcal{X}}_{\sigma})$, and $N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{\beta}(\mathcal{X}_{\sigma})$ for every $\beta\in NE(\mathcal{X}_{\sigma}/\overline{\mathcal{X}}_{\sigma})$.
@@
Let $Z_{r}$ be the stratum of $\mathcal{X}_{0}$ corresponding to the smallest cone of $C(P)$ containing $r$. As $r\in C(\sigma)_{\mathbb{Z}}$, it follows from Lemma 7.8 that $x_{\sigma}\in f_{H}(Z_{r})$. Recall that the Definition 5.3 of $N_{pqr}^{\beta}(\mathcal{X})$ involves considering punctured log curves passing through a given point $z$ in the interior of $Z_{r}$. As $U_{\sigma}$ is an open neighborhood of $x_{\sigma}$, there exists a point $z$ in the interior of $Z_{r}$ such that $f_{H}(z)\in U_{\sigma}$. If one uses such a point $z$ to calculate $N_{pqr}^{\beta}(\mathcal{X})$, then all the contributing punctured log curves are contained in $f_{H}^{-1}(z)$ and so in $\mathcal{X}_{\sigma,0}$, since $\beta\in H$ and so curves of class $\beta$ are contracted by $f_{H}$. In particular, we necessarily have $\beta\in NE(\mathcal{X}_{\sigma}/\overline{\mathcal{X}}_{\sigma})$ if $N_{pqr}^{\beta}(\mathcal{X})\neq 0$. Finally, as all these curves are contained in $\mathcal{X}_{\sigma,0}$, and the log structure on $\mathcal{X}_{\sigma,0}$ is the restriction of the log structure on $\mathcal{X}_{0}$, we have $N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{\beta}(\mathcal{X}_{\sigma})$ by Lemma 5.6. ∎
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration and let $H$ be a face of $NE(\mathcal{X}/\overline{\mathcal{X}})$. Then, the following holds:
@@
    For every 2-dimensional cell $\sigma$ of $\mathscr{P}_{H}$, denoting by $\mathcal{Y}_{\sigma}$ the corresponding irreducible component of $\mathcal{Y}_{\mathcal{X}}|_{S_{H}}$, by $\partial\mathcal{Y}_{\sigma}$, $\mathcal{D}_{\sigma}$ and $\mathcal{C}_{\sigma}$ the restrictions of the double locus of $\mathcal{Y}_{\mathcal{X}}|_{S_{H}}$, of $\mathcal{D}_{\mathcal{X}}|_{S_{\mathcal{X}}}$ and $\mathcal{C}_{\mathcal{X}}|_{S_{\mathcal{X}}}$ to $\mathcal{Y}_{\sigma}$, the family $(\mathcal{Y}_{\sigma},\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}+\epsilon\,\mathcal{C}_{\sigma})\rightarrow S_{H}$ is the base change, along the natural map $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$ as in Equation (7.1), of the mirror family $(\mathcal{Y}_{\mathcal{X}_{\sigma}},\mathcal{D}_{\mathcal{X}_{\sigma}}+\epsilon\,\mathcal{C}_{\mathcal{X}_{\sigma}})$ to the quasi-projective open Kulikov degeneration $\mathcal{X}_{\sigma}\rightarrow\Delta$.
@@
Denoting $J_{H}:=NE(\mathcal{X}/\overline{\mathcal{X}})\setminus H$, we have $\mathcal{Y}_{\mathcal{X}}|_{S_{H}}=\mathrm{Proj}\,\mathcal{R}_{\mathcal{X}}/J_{H}$. Explicitly, $\mathcal{R}_{\mathcal{X}}/J_{H}$ is the algebra of theta functions $(\vartheta_{p})_{p\in C_{\mathbb{Z}}}$ where the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ with $\beta\notin H$ are set to zero. By Lemma 7.10, the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ with $\beta\in H$ are zero unless $p,q,r$ are contained in a common cone $C(\sigma)$ over a 2-dimensional cell $\sigma$ of $\mathscr{P}_{H}$. Therefore, for every such $\sigma$, denoting by $\mathcal{R}_{\sigma}$ the $\mathbb{C}[H]$-algebra spanned by the theta functions $(\vartheta_{p})_{p\in C(\sigma)_{\mathbb{Z}}}$, it follows that $\mathcal{Y}_{\sigma}:=\mathrm{Proj}\,\mathcal{R}_{\sigma}$ is an union of irreducible components of $\mathcal{Y}_{\mathcal{X}}|_{S_{H}}$. Moreover, denoting by $\partial\mathcal{Y}_{\sigma}$ the union of $\mathcal{Y}_{\sigma}\cap\mathcal{Y}_{\sigma^{\prime}}$ with $\sigma^{\prime}\neq\sigma$, it follows that $\partial\mathcal{Y}_{\sigma}$ is defined by $\vartheta_{p}=0$ for every $p\in C(\sigma)_{\mathbb{Z}}$ such that there does not exist $\sigma^{\prime}\neq\sigma$ with $p\in C(\sigma^{\prime})$. On the other hand, by the definition of $\mathcal{D}_{\mathcal{X}}$ in Equations (6.8)-(6.9), the restriction $\mathcal{D}_{\sigma}:=\mathcal{D}_{\mathcal{X}}|_{\mathcal{Y}_{\sigma}}$ is defined by $\vartheta_{p}=0$ for $p\notin(C(\partial\sigma)\cap C(\partial P))_{\mathbb{Z}}$. For every $p\in C(\partial\sigma)_{\mathbb{Z}}$, we have $p\in C(\sigma^{\prime})$ for some $\sigma^{\prime}\neq\sigma$ or $p\in(C(\partial\sigma)\cap C(\partial P))_{\mathbb{Z}}$, and so the divisor $\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}$ is defined by $\vartheta_{p}=0$ for $p\in C(\sigma)_{\mathbb{Z}}\setminus C(\partial\sigma)_{\mathbb{Z}}$.

### v3

@@ 7.1. Generic KSBA stability of the mirror family
To prove ii), it suffices to prove that $Y_{t}$ is normal, since then the other conditions follow by i). By Theorem 6.36, the pair $(Y_{t},D_{t})$ is semi-log-canonical. In particular, the surface $Y_{t}$ is $S_{2}$, and so, by Serre’s criterion for normality [56, Thm 8.22A], it suffices to show that $Y_{t}$ is regular in codimension one. As in the proofs of Lemma 6.29 and Theorem 6.36, we can assume that there exists a good divisor $F$ on $\mathcal{X}$, inducing a $\mathbb{C}^{\star}$-action on $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$. The closure of the $\mathbb{C}^{\star}$-orbit passing through $t$ is an affine line $\mathbb{A}^{1}$, with $(\mathbb{A}^{1}\setminus\{0\})\subset T_{\mathcal{X}}$, and $0\in\mathrm{Spec}\,\mathbb{C}[G]\simeq\mathrm{Spec}\,\mathbb{C}[H]/J$, where $G=F^{\perp}\cap NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ and $J=NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})\setminus G$. The restriction $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$ of $\mathcal{Y}_{\mathcal{X}}$ to this affine line is constant with fibers $Y_{t}$ on $\mathbb{A}^{1}\setminus\{0\}$. Moreover, by the proof of Lemma 6.28, the central fiber $\mathcal{Y}_{\mathcal{X},0}$ of $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$ is a union of toric varieties with intersection complex $(P,\mathscr{P}_{J})$, where $\mathscr{P}_{J}$ is the subdivision of $P$ obtained from $\mathscr{P}_{J}$ by omitting the edges $e$ such that the corresponding curves $C_{e}$ in the double locus of $\mathcal{X}_{0}$ satisfy $F\cdot C_{e}=0$. In particular, the singular locus of $(P,\mathscr{P}_{J})$ is the union of its double curves $D_{e}$, which are in one-to-one correspondence with the edges $e$ of $\mathscr{P}_{J}$, that is, with the double curves $C_{e}$ of $\mathcal{X}_{0}$ such that $F\cdot C_{e}\neq 0$. Hence to prove that $Y_{t}$ is smooth in codimension one, it suffices to show that each of the double curves $D_{e}$ are generically smoothed out in $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$.
@@
It follows from the construction of $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ from the canonical scattering diagram, together with [48, Eq.(2.17)], that a neighborhood of $D_{e}$ in $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$ is isomorphic to the affine variety of equation $xy=t^{[C_{e}]}f_{W,e}$, where $f_{W,e}$ is the function of the canonical scattering diagram attached to the wall supported on the edge $e$ of $P$. As $(\mathbb{A}^{1}\setminus\{0\})$ is contained in $T_{\mathcal{X}}$, we have $t^{[C_{e}]}\neq 0$ on the non-central fibers of $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$. Moreover, by Lemma 6.17, the function $f_{W,e}$ is generically non-zero on $D_{e}$. This shows that all the double curves $D_{e}$ of $\mathcal{Y}_{\mathcal{X},0}$ are generically smoothed out in the one-parameter family $\mathcal{Y}_{\mathcal{X}}|_{\mathbb{A}^{1}}$, and this concludes the proof that $Y_{t}$ is normal. ∎
@@ 7.2. KSBA stability of the mirror family
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration, and let $H$ be a face of $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$. As $f:\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ is a relative Mori dream space by Proposition 3.36, there is a corresponding birational contraction $f_{H}:\mathcal{X}\rightarrow\mathcal{X}_{H}$ such that $H=NE(\mathcal{X}/\mathcal{X}_{H})$. Denote by $S_{H}:=\mathrm{Spec}\,\mathbb{C}[H]=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\mathcal{X}_{H})]$ the corresponding stratum of $S_{\mathcal{X}}$, with dense torus orbit $T_{H}:=\mathrm{Spec}\,\mathbb{C}[H^{\mathrm{gp}}]$. Let $\mathscr{P}_{H}$ be the polyhedral decomposition of $P$ obtained as a coarsening of $\mathscr{P}$ by removing all edges $e$ of $\mathscr{P}$ such that the corresponding double curves $C_{e}$ of $\mathcal{X}_{0}$ satisfy $C_{e}]\in H$, that is, are contracted by $f_{H}$. For every 2-dimensional cell $\sigma$ of $\mathscr{P}_{H}$, we denote by $\mathscr{P}_{\sigma}$ the polyhedral decomposition of $\sigma$ obtained by restricting $\mathscr{P}$ to $\sigma$.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration and let $H$ be a face of $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$. Then, for every 2-dimensional cell $\sigma$ of $\mathscr{P}_{H}$, there exists a generic quasi-projective open Kulikov surface $\mathcal{X}^{\prime}_{\sigma,0}$, with dual intersection complex $(\sigma,\mathscr{P}_{\sigma})$, such that $\mathcal{X}_{\sigma,0}$ is a Zariski open dense subset of $\mathcal{X}^{\prime}_{\sigma,0}$, containing all compact curves of $\mathcal{X}^{\prime}_{\sigma,0}$, and $U_{\sigma}$ is a Zariski open neighborhood of the central point in the affinization $(\mathcal{X}^{\prime}_{\sigma,0})^{\mathrm{can}}$.
@@
First note that $\mathcal{X}_{\sigma,0}$ is quasi-projective, reduced and normal crossings, as an open subset of the quasi-projective, reduced and normal crossings surface $\mathcal{X}_{0}$. If $v\in\sigma_{\mathbb{Z}}$ and $v\notin\partial\sigma$, then, by Lemma 7.2 ii), we have $X^{v}_{\sigma}=X^{v}$, and so $X^{v}_{\sigma}$ is a 2-surface. If $v\in\partial\sigma$, then it follows from Definitions 3.2-3.4 that $X^{v}_{\sigma}$ is naturally a Zariski open subset of a 0-surface or a 1-surface if $\dim f_{H}(X^{v}_{\sigma})=2$ or $\dim f_{H}(X^{v}_{\sigma})=1$ respectively. Finally, it follows from [90, Proposition 3.3] that $\sigma$ is homeomorphic to a disk. ∎
Let $\mathcal{X}^{\prime}_{\sigma,0}$ be the quasi-projective open Kulikov surface given by Lemma 7.3. By Theorem 3.31, we can assume up to locally trivial deformations that $\mathcal{X}_{\sigma,0}^{\prime}$ is generic $d$-semistable, and so that $\mathcal{X}_{\sigma,0}^{\prime}$ is the central fiber of a quasi-projective open Kulikov degeneration $\mathcal{X}_{\sigma}\rightarrow\Delta$ by Theorem 3.19 and Lemma 3.24. The inclusions $\mathcal{X}_{\sigma,0}\subset\mathcal{X}_{0}$ and $\mathcal{X}^{\mathrm{can}}_{\sigma,0}\subset\mathcal{X}_{H,0}$ induce a map $NE(\mathcal{X}_{\sigma}/\mathcal{X}^{\mathrm{can}}_{\sigma})\rightarrow H=NE(\mathcal{X}/\mathcal{X}_{H})$, and so a toric morphism
$S_{H}=\mathrm{Spec}\,\mathbb{C}[H]\longrightarrow S_{\mathcal{X}_{\sigma}}=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\sigma}/\mathcal{X}^{\mathrm{can}}_{\sigma})]\,.$ (7.1)
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration, and let $H$ be a face of $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$. Then, for every $\beta\in H$ and $p,q,r\in C(P)_{\mathbb{Z}}$, we have $N_{pqr}^{\beta}(\mathcal{X})=0$ unless $p,q,r$ are all contained in a common cone $C(\sigma)$ over a 2-dimensional cell $\sigma$ of $\mathscr{P}_{H}$.
@@
The tropicalization of a punctured log curve $C$ contributing to $N_{pqr}^{\beta}(\mathcal{X})$ as in Definition 5.3 is a connected tropical curve in $C(P)$, with two unbounded edges with asymptotic directions $\mathbb{R}_{\geq 0}p$ and $\mathbb{R}_{\geq 0}q$, and with one leg contained in $\mathbb{R}_{\geq 0}r$. In particular, if $p,q,r$ were not all contained in a common cone $C(\sigma)$, this tropical curve would have to intersect a cone $C(\sigma\cap\sigma^{\prime})$ over the intersection of two distinct 2-dimensional faces $\sigma$, $\sigma^{\prime}$ of $\mathscr{P}_{\sigma}$, and so a cone $C(e)$ over an edge $e$ of $\mathscr{P}_{\sigma}$. In particular, the curve $C$ would contain an irreducible component with class a multiple of $C_{e}]$, where $C_{e}$ is the double curve of $\mathcal{X}_{0}$ corresponding to $e$. Since $e$ is an edge of $\mathscr{P}_{\sigma}$, we have $C_{e}]\notin H$, and this would imply $\beta\notin H$. Hence, the results follows since we assume that $\beta\in H$. ∎
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration and let $H$ be a face of $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$. Then, for every 2-dimensional cell $\sigma$ of $\mathscr{P}_{H}$, for every $p,q,r\in C(\sigma)_{\mathbb{Z}}$ and $\beta\in H$, we have $N_{pqr}^{\beta}(\mathcal{X})=0$ if $\beta\notin NE(\mathcal{X}_{\sigma}/\mathcal{X}^{\mathrm{can}}_{\sigma})$, and $N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{\beta}(\mathcal{X}_{\sigma})$ for every $\beta\in NE(\mathcal{X}_{\sigma}/\mathcal{X}^{\mathrm{can}}_{\sigma})$.
@@
Let $Z_{r}$ be the stratum of $\mathcal{X}_{0}$ corresponding to the smallest cone of $C(P)$ containing $r$. As $r\in C(\sigma)_{\mathbb{Z}}$, it follows from Lemma 7.2 that $x_{\sigma}\in f_{H}(Z_{r})$. Recall that the Definition 5.3 of $N_{pqr}^{\beta}(\mathcal{X})$ involves considering punctured log curves passing through a given point $z$ in $Z_{r}$. As $U_{\sigma}$ is an open neighborhood of $x_{\sigma}$, there exists a point $z$ in $Z_{r}$ such that $f_{H}(z)\in U_{\sigma}$. Using such a point $z$ to calculate $N_{pqr}^{\beta}(\mathcal{X})$, all the contributing punctured log curves are contained in $f_{H}^{-1}(f_{H}(z))$ and so in $\mathcal{X}_{\sigma,0}$, since $\beta\in H$ and so curves of class $\beta$ are contracted by $f_{H}$. In particular, we necessarily have $\beta\in NE(\mathcal{X}_{\sigma}/\mathcal{X}^{\mathrm{can}}_{\sigma})$ if $N_{pqr}^{\beta}(\mathcal{X})\neq 0$. Finally, as all these curves are contained in $\mathcal{X}_{\sigma,0}$, and the log structure on $\mathcal{X}_{\sigma,0}$ is the restriction of the log structure on $\mathcal{X}_{0}$, we have $N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{\beta}(\mathcal{X}_{\sigma})$ by Lemma 5.7. ∎
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration and let $H$ be a face of $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$. Then, the following holds:
@@
    For every 2-dimensional cell $\sigma$ of $\mathscr{P}_{H}$, denote by $\mathcal{Y}_{\sigma}$ the corresponding irreducible component of $\mathcal{Y}_{\mathcal{X}}|_{S_{H}}$, and by $\partial\mathcal{Y}_{\sigma}$, $\mathcal{D}_{\sigma}$ and $\mathcal{C}_{\sigma}$ the restrictions to $\mathcal{Y}_{\sigma}$ of the double locus of $\mathcal{Y}_{\mathcal{X}}|_{S_{H}}$, $\mathcal{D}_{\mathcal{X}}|_{S_{\mathcal{X}}}$ and $\mathcal{C}_{\mathcal{X}}|_{S_{\mathcal{X}}}$ respectively. Then, the family $(\mathcal{Y}_{\sigma},\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}+\epsilon\,\mathcal{C}_{\sigma})\rightarrow S_{H}$ is the base change along $S_{H}\rightarrow S_{\mathcal{X}_{\sigma}}$ of the mirror family $(\mathcal{Y}_{\mathcal{X}_{\sigma}},\mathcal{D}_{\mathcal{X}_{\sigma}}+\epsilon\,\mathcal{C}_{\mathcal{X}_{\sigma}})$ to the quasi-projective open Kulikov degeneration $\mathcal{X}_{\sigma}\rightarrow\Delta$.
@@
Denoting $J_{H}:=NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})\setminus H$, we have $\mathcal{Y}_{\mathcal{X}}|_{S_{H}}=\mathrm{Proj}\,\mathcal{R}_{\mathcal{X}}/J_{H}$. Explicitly, $\mathcal{R}_{\mathcal{X}}/J_{H}$ is the algebra of theta functions $(\vartheta_{p})_{p\in C_{\mathbb{Z}}}$ where the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ with $\beta\notin H$ are set to zero. By Lemma 7.4, the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ with $\beta\in H$ are zero unless $p,q,r$ are contained in a common cone $C(\sigma)$ over a 2-dimensional cell $\sigma$ of $\mathscr{P}_{H}$. Therefore, for every such $\sigma$, denoting by $\mathcal{R}_{\sigma}$ the $\mathbb{C}[H]$-algebra spanned by the theta functions $(\vartheta_{p})_{p\in C(\sigma)_{\mathbb{Z}}}$, it follows that $\mathcal{Y}_{\sigma}:=\mathrm{Proj}\,\mathcal{R}_{\sigma}$ is an union of irreducible components of $\mathcal{Y}_{\mathcal{X}}|_{S_{H}}$. Moreover, denoting by $\partial\mathcal{Y}_{\sigma}$ the union of $\mathcal{Y}_{\sigma}\cap\mathcal{Y}_{\sigma^{\prime}}$ with $\sigma^{\prime}\neq\sigma$, it follows that $\partial\mathcal{Y}_{\sigma}$ is defined by $\vartheta_{p}=0$ for every $p\in C(\sigma)_{\mathbb{Z}}$ such that there does not exist $\sigma^{\prime}\neq\sigma$ with $p\in C(\sigma^{\prime})$. On the other hand, by the definition of $\mathcal{D}_{\mathcal{X}}$ in Equations (6.10)-(6.11), the restriction $\mathcal{D}_{\sigma}:=\mathcal{D}_{\mathcal{X}}|_{\mathcal{Y}_{\sigma}}$ is defined by $\vartheta_{p}=0$ for $p\notin(C(\partial\sigma)\cap C(\partial P))_{\mathbb{Z}}$. For every $p\in C(\partial\sigma)_{\mathbb{Z}}$, we have $p\in C(\sigma^{\prime})$ for some $\sigma^{\prime}\neq\sigma$ or $p\in(C(\partial\sigma)\cap C(\partial P))_{\mathbb{Z}}$, and so the divisor $\mathcal{D}_{\sigma}+\partial\mathcal{Y}_{\sigma}$ is defined by $\vartheta_{p}=0$ for $p\in C(\sigma)_{\mathbb{Z}}\setminus C(\partial\sigma)_{\mathbb{Z}}$.